40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Engineering group Lab elog, Page 3 of 6  Not logged in ELOG logo
ID Date Author Type Category Subject
  164   Tue Jul 16 14:54:47 2019 Andy R.ProgressBS BRDsBRD3 Calibration

BRD3_v4 tuned on 7/15 with the following data: (f and Q averaged over 4 measurements)

  Bounce Roll
Mass

2.352g+.946g+.949g+screw=4.865g

2.625g+screw+spacer=3.207g
f 16.648 Hz 24.348 Hz
Q 153.68 95.58
Measurement #s 31-34 25-28

 

BRD3_v4 Drifts to 7/16, masses unchanged: (averaged over 4 measurements):

  Bounce Roll
f 16.692 Hz 24.416 Hz
Q 158.48 101.23
Measurement #s 35-38 39-42

So the net drift overnight is:

  • +0.044 Hz for bounce ( i.e. +0.3%).  Mistuning is now 0.01%. 
  • +0.068 Hz for roll (i.e. +0.3%). Mistuning is now 0.3%.
  163   Fri Jul 12 12:03:54 2019 Marie K.ProgressBS BRDsFurther measurements of BS with BRDs

Checking the BS modes this morning (July 12), we observed a change in the bounce mode behaviour, which we assume is due to a shift in BRD2_v4 resonance (tuning described in elog 160). We now have a split mode!

Bounce

B1 Freq. [Hz]

Q1 (from 3dB width) B2 Freq. [Hz] Q2 (from 3dB width) Measured range  Nb points Drive [V] Filename Comments
Measurement B5 16.637   16.730   16.5 - 17 600 0.3 srs0011.txt BRD2_v4 on

 

Mesuring the roll mode again, we observed a shift in the mode frequency as well. However, this doesn't change the Q of the modes.

Roll

B1 Freq. [Hz]

Q1 (from 3dB width) B2 Freq. [Hz] Q2 (from 3dB width) Measured range  Nb points Drive [V] Filename Comments
Measurement R9 24.15 193 24.476 215 24 - 24.6 600 0.01 srs0012.txt BRD2_v4 on
Measurement R10 24.13 155 24.448 247 24 - 24.6 600 0.1 srs0013.txt BRD2_v4 on
Attachment 1: BRD2_R9_R10.png
BRD2_R9_R10.png
Attachment 2: BRD2_B5.png
BRD2_B5.png
  162   Fri Jul 12 11:04:02 2019 Andy R.ProgressBS BRDsDrift and Retuning of BRD1

Overnight, the BRD1 drifted to resonant frequencies (paranthesis are net drift):

Bounce: 16.792 Hz (+.112 Hz)

Roll: 24.496 Hz (+.76 Hz)

BRD1 was then retuned to the following values as a mean over 4 measurements:

Bounce: f=16.64 Hz, Q=161.74 (Measurements 43-46)

Roll: f=24.37 Hz, Q=86.5 (Measurements 35-38)

  161   Thu Jul 11 16:16:32 2019 Marie K.ProgressBS BRDsBRD2v4 installed on BS

Today we installed BRD2_v4 (tuned described in elog 160) on the left side of BS. Scanning over a large frequency range (as in elog 152 - 14 to 26 Hz ), we observed a nice split of the roll mode (see figure 1 - saved as SRS007.78D).

Scanning over shorter frequency spans, we measured the two modes:

Roll

R1 Freq. [Hz]

Q1 (from 3dB width) R2 Freq. [Hz] Q2 (from 3dB width) Measured range  Nb points Drive [V] Filename
Measurement R7 24.119 +/- 0.002 147 24.447 +/- 0.002 248 24 - 25 600 0.01 SRS008.78D
Measurement R8 24.121 +/- 0.001 147 24.447 +/- 0.001 254 24.0 - 24.6 600 0.01 SRS009.78D

 

The bounce mode split is not as easy to observe, as it is reported in T1700176.

Bounce

B1 Freq. [Hz]

Q1 (from 3dB width) B2 Freq. [Hz] Q2 (from 3dB width) Measured range  Nb points Drive [V] Filename Comments
Measurement B3 16.644 +/- 0.003 300 - - 16-18 600 0.01   No split observed
Measurement B4         16.5 - 16.7     SRS010.78D No split observed (see picture)

Tomorrow we will add BRD1_v4 to the BS suspension and remeasure the modes.

Note: I am somewhat dubious concerning the state of the blade on the bounce side, because I torqued it too much during a tuning and it has a permanent twist: we can see a remaining bend on the blend under the right light. We should change the blade and retune the BRD.

 

Attachment 1: BS_withBRD2_v4.jpg
BS_withBRD2_v4.jpg
Attachment 2: MeasR7.jpg
MeasR7.jpg
Attachment 3: MeasR8.jpg
MeasR8.jpg
Attachment 4: MeasB4.jpg
MeasB4.jpg
  160   Thu Jul 11 11:20:21 2019 Andy R.ProgressBS BRDsBRD2 Calibration

BRD2 modes right before being tested in suspension as a mean of 4 measurements:

Bounce: f=16.678 Hz, Q=146.042  (Measurements 129-132)

Roll: f=24.236 Hz, Q=99.02  (Measurements 133-136)

 

BRD1 modes as of noon as a mean of 4 measurements:

Bounce: f=16.68 Hz, Q=200.91 (Measurements 13-16)

Roll: f=24.42 Hz, Q=96.6 (Measurements 17-20)

  159   Thu Jul 11 10:41:22 2019 Luis SanchezElectronicsA-plusHSTS update wiring and bom diagrams

Luis:

An update from A+ HSTS documents were completed this include block diagram document E1900106, wiring diagran document D1900169 and BOM document E1900171. Also I created a collector point for the A+ HSTS review, the document is E1900183.

In E1900183 document , HAM7 bill of materials is seen(also in the attached files below) this document include the Tip-Tilts and HSTS structures. TIP-TILT material is considering 3 original TIP-TILT units and 2 HDS, the HDS include 4 AOSEMs and 4 Coils.

 

Attachment 1: Aplus_HSTS_Electronics_Wiring.pdf
Aplus_HSTS_Electronics_Wiring.pdf Aplus_HSTS_Electronics_Wiring.pdf Aplus_HSTS_Electronics_Wiring.pdf Aplus_HSTS_Electronics_Wiring.pdf Aplus_HSTS_Electronics_Wiring.pdf Aplus_HSTS_Electronics_Wiring.pdf Aplus_HSTS_Electronics_Wiring.pdf Aplus_HSTS_Electronics_Wiring.pdf
Attachment 2: Electronics_BOM_HAM7(HSTS-TIP_TILT).pdf
Electronics_BOM_HAM7(HSTS-TIP_TILT).pdf
  Draft   Thu Jul 11 10:29:02 2019 Luis SanchezElectronicsA-plusHSTS update wiring and bom diagrams

Luis:

An update from A+ HSTS documents were completed this include block diagram document E1900106, wiring diagran document D1900169 and BOM document E1900171. Also I created a collector point for the A+ HSTS review, the document is E1900183.

In E1900183 document , HAM7 bill of materials is seen(also in the attached files below) this document include the Tip-Tilts and HSTS structures. TIP-TILT material is considering 3 original TIP-TILT units and 2 HDS, the HDS include 4 AOSEMs and 4 Coils.

 

  157   Wed Jul 10 15:03:25 2019 Marie K.ProgressBS BRDsRing down of the BS modes

Since we know the bounce and roll mode frequencies well, I decided today to measure the Q of the in-air dummy BS suspension by fitting the ring-down of the modes. This will help us quantify the damping effect of the BRDs when added on the suspension.

The fits are computed over the envelop of the decay (see attached code). I added for reference the tau of the fits, which can to quantify the decay without asomption on the resonance frequency :

decay = A0 exp(-t/tau) = A0 exp(-wt/(2Q))  = A0 exp(-2*pi*f*t/(2Q))

  • For the bounce mode, the mean Q over 9 mesurements is 4614 (see figure 1 attached)
  • For the roll mode, the mean Q over 7 measurements is 4506 (see figure attached, ignoring the fits #5 and #9 that are poor measurements).

In both cases, the Q meaured with this method are higher than what we measure with the 3dB method.

Attachment 1: BS_bounce_ringdown_noBRD_July10.png
BS_bounce_ringdown_noBRD_July10.png
Attachment 2: BS_roll_ringdown_noBRD_July10.png
BS_roll_ringdown_noBRD_July10.png
Attachment 3: BS_ring_down_meas_DAQ.m
%% BRDs Ring Down Measurements
%  - Read the files
%  - Estimate the Q from the decay fit
%
% MK May 2019

clc; clear all; close all;

%% Read the data from .csv file
data_dir = 'C:\Users\mkasprza\Documents\SUS_BRDs\BS_tests\';
... 87 more lines ...
  156   Wed Jul 10 11:56:10 2019 Andy R.ProgressBS BRDsTuning of BRD2

Tuning of BRD2 modes and both are in the target range

Roll Mode (mean of 3 measurements):  f=24.240 Hz,  Q=100.34

Bounce Mode (mean of 4 measurements): f=16.708 Hz,  Q=134.88

  155   Tue Jul 2 09:00:46 2019 Marie K. ProgressBS BRDsBRD2 tuning

Retuning BRD2 after yesterday trials. After moving the masses again along the blade axis, I get (mean over 6 measurements):

  • bounce mode: f = 16.76, Q = 130, tuning = 0.4%
  • roll mode: f = 24.38, Q = 126, tuning = 0.2%
  154   Fri Jun 28 16:05:40 2019 Marie K. ProgressBS BRDsBRD2 tuning

I tuned BRD2 (with version 4 of the blades) to the measured frequencies of the BS, aiming for a tuning better than 0.5 % as displayed in the range below. 

After some trials, I got the bounce mode tuned to 0.2% and the roll mode tuned better than 0.1 %.

Masses/Freq

Bounce - target = 16.69 Hz

(range = 16.61-16.77 Hz)

Resonance [Hz]

Roll - target = 24.34 Hz

(range = 24.22 - 24.46 Hz)

Resonance [Hz]
Test 1 2.385+1.029+1.026+screw = 5.050 g 15.450 2.546+screw+spacer = 3.198g 25.175
Test 2 Masses moved along the axis to raise the freq. Max position. 16.400 Masses moved along the axis to lower the freq. Max position. 23.075
Test 3

less mass is needed & moved to max position

2.383+1.025+0.987+screw = 5.006 g

16.800 Masses moved back and forth few times 24.325
Test 4 Masses moved back and forth a few times - f still too low      
Test 5 2.383+2*0.98+screw = 4.967g - still too low 16.475    
Test 6 2.383+0.98+0.95+screw = 4.927 g 16.575    
Test 7 2.383+0.95*2+screw = 4.889 g + back and forth tuning 16.700 - 16.725    

However, repeating the measurements several times after the tuning, I got the following results (mean over 6 measurements):

  • bounce mode: f = 16.750 Hz, Q = 131, tuning = 0.4% (see figure 1)
  • roll mode: f = 24.496 Hz, Q = 133, tuning = 0.6% (see figure 2)

It turns out the tuning is not that great. Furthermore, the Q are higher than expected from previous tests. When measuring the decay, the values were closer to 90. I should rerun the Q analysis via decay fit over the same set of measurements.

Attachment 1: brd2v4_bounce_tuned_july12019.png
brd2v4_bounce_tuned_july12019.png
Attachment 2: brd2v4_roll_tuned_july12019.png
brd2v4_roll_tuned_july12019.png
  153   Fri Jun 28 08:15:41 2019 Marie K.ProgressBS BRDsBS resonances

Starting to measure the BS resonances more finely.

Bounce Freq [Hz] Q (from 3dB width) Measured range Nb points Drive [V] Filename Comments
Measurement B1 16.6937 +/- 0.0008 909 16.5 - 17.0 600 0.1   no BRD
Measurement B2 16.6922 +/- 0.0001 2200 16.675-16.710 600 0.1   no BRD

The measured Q is low for the first measurement, but might be due to the low resolution. Measurement B2 is closer to what is expected from previous measurements (see Norna's document T1700176-v7, page 4): bounce mode frequency = 16.70 Hz, Q = 3400. One or two more measurements would be useful to confirm the Q.

 

Roll Freq. [Hz] Q (from 3dB width) Measured range  Nb points Drive [V] Filename Comments
Measurement R1 24.341 +/- 0.003 - 24 - 25 600     no BRD
Measurement R2 24.33477 +/- 0.00005 2960 24.320 - 24.350 600 0.01 SRS001.78D no BRD
Measurement R3 24.33722 +/- 0.00003 3320 24.330 - 24.345 600 0.01 SRS003.78D no BRD
Measurement R4 24.33601 +/ 0.00003 3294 24.330 - 24.345 600 0.05 SRS004.78D? no BRD
Measurement R5 24.33441 +/- 0.00003 2500 24.330 - 24.345 600 0.03 SRS005.78D no BRD
Measurement R6 24.33619 +/- 0.00003 2884 24.330 - 24.345 600 0.01 SRS006.78D no BRD

To be compared to Norna's results: Roll mode frequency = 24.34 Hz, Q = 3000.

Attachment 1: MeasB1.jpg
MeasB1.jpg
Attachment 2: MeasB2.jpg
MeasB2.jpg
Attachment 3: MeasR1.jpg
MeasR1.jpg
Attachment 4: MeasR2.jpg
MeasR2.jpg
Attachment 5: MeasR3.jpg
MeasR3.jpg
  152   Wed Jun 26 15:18:44 2019 Marie K.ProgressBS BRDsMeasurement of the resonance frequencies - 1

Today I was able to measure the BS bounce and roll frequencies (see figure attached). The previous measurements were not working because channel 1 of the signal analyzer was not plugged in to the source (thank you Rich for pointing this out!). Rich explained to me that the transfer functions are by design always CH2/CH1. The 3 BNC connections of the analyzer are independent. In our case, to know the BOSEM response over the excitation provided by the analyzer, one needs to connect CH1 directly to the output of the source and CH2 to the readout of the bosem. The measurements are obviously now much cleaner and ressemble what Norna had measured in 2017.

  • Measurement of the dummy BS resonances:
    • Precision of 0.024 Hz (2.77 ks measurement - 500 pts from 14 Hz to 26 Hz, with 5s int. time and 5 int. cycle per points). Sine wave excitation of 0.1V. There was no saturation of the readout at the max. of the peaks.
    • Bounce mode frequency: 16.74 Hz
    • Roll mode frequency: 24.39 Hz
    • More measurements are now needed around the resonances to determine the Q of each mode.

I also started to remeasure the BRDs in order to tune them, this time with an excitation to clearly see the resonance frequency.

  • Measurements of the stand-alone BRDs with manual excitation:
    • Precision of 0.03 Hz (32s measurement) (~0.2% for bounce, ~0.1% for roll)
    • With an average over 5 measurements for each mass, the BRDs frequencies are:
      • BRD2 bounce = 16.88 Hz (Q~16.88/0.134 = 126) - roll = 24.906 Hz
    • We now have to tune the masses to match the BS resonances measured as explained below.
Attachment 1: BS_TF.jpg
BS_TF.jpg
  151   Mon Jun 24 17:19:24 2019 Marie K.ProgressBS BRDsHigher resolution spectra

We need to tune the BRDs at the right frequency with a precision better than 0.5%. It means we need a to know the frequency better than 0.08 Hz for the bounce mode (0.1 Hz for roll). Thanks to Stephen who kindly explained me the B&K software today, I remeasured the bounce and roll frequency of the already assembled BRDs with a greater precision.

  • BRD1 frequencies: 17.100 Hz  -  24.300 Hz
  • BRD2 frequencies: 16.750 Hz  - 24.913 Hz

However the peaks are too broad to really state what is the actual frequency, so these numbers aren't very meaningful for now. I should retake measurements with an excitation on the BRDs. There is a sharp peak at 23.275 Hz that is present on all measurements.

For the record: the resolution of the measurement can be set under the tab "Measure" --> "Standard Measurement"--> "Measurement Mode"and then in "FFT system" we can define the Spectrum Parameters (see screenshot).

The FFT parameters that are used are 10 avg - avg time 320s - FFt Line 800 - Freq resolution 0.0125

Attachment 1: resultbrd_screenshot.png
resultbrd_screenshot.png
Attachment 2: setup_fft_brd_screenshot.png
setup_fft_brd_screenshot.png
  150   Fri Jun 21 12:57:38 2019 Marie K.ProgressBS BRDsTheoretical best achievable Q as a function of mass ratio

Posting here for reference the best achievable Q for the damped BS as a function of the mass ratio of the damper mass and BS mass (see figure attached, along with the Matlab code). This is mainly based on a code from Norna (T1600259), with a line that computes the structural damping constant of the damper from the mass ratio (see Equation 12 in T1500271).

Currently we are using masses of 5g for bounce (3g for roll) for the damper. The modal masses of the BS are 27.78 kg (bounce mode) and 12.9 kg (roll mode). The mass ratio is therefore 1.8e-4 (2.3e-4). It means that we could achieve a Q ~ 120 in case of perfect tuning or Q ~ 200 for a mistuning by 0.5% for the bounce mode.

For the roll mode, we could do slightly better, with a Q~ 110 if the damper is perfectly tuned, and a Q ~170 for 0.5% detuning.

Attachment 1: Best_Q_vs_mistuning.png
Best_Q_vs_mistuning.png
Attachment 2: BRD_HSTS_T1600259_loop_for_graph.m
% Create a loop around Norna's model to reproduce Brett's graph
% from T1500271 + adjust the damping 
%
% Model from Norna concerning the BRD_HSTS
% parameters adjusted for the BS dampers v4
% MK May 2019

%model of coupled oscillator system: blade and damper
%damping in damper only
% NAR 5th May 2015
... 98 more lines ...
  149   Tue Jun 11 07:58:39 2019 Luis SanchezElectronicsA-plusSome other Wiring Diagrams and Block Diagrams related to Aplus

Luis:

I had been creating some more block and wiring diagrams on visio and altium. These diagrams are related to the Readout (E1900165, D1900190) and Original Tip Tilt (E1900163, D1900191) suspension assemblys. The concept of tip tilt suspension is still in process of develop and will have some changes, they need to define which will required some Active Mode Matching control and/or  HDS with SAMs control. 

I added a sheet on Readout block diagram where we use a single Sat Box (with 8ch instead 2 Sat Box with 4ch) document E1900165. Also,  for this diagram I added a HAM A Coil driver block.

I also added on DCC a Bill of Materials for HSTS E1900171, HRTS E1900172.

Some of this diagrams can be seen in attached files.

Attachment 1: Suspension_Wiring_Block_Diagram_Readout_Platform_Suspension_E1900165-v2.pdf
Suspension_Wiring_Block_Diagram_Readout_Platform_Suspension_E1900165-v2.pdf Suspension_Wiring_Block_Diagram_Readout_Platform_Suspension_E1900165-v2.pdf Suspension_Wiring_Block_Diagram_Readout_Platform_Suspension_E1900165-v2.pdf Suspension_Wiring_Block_Diagram_Readout_Platform_Suspension_E1900165-v2.pdf
Attachment 2: BOM_List_HSTS.pdf
BOM_List_HSTS.pdf BOM_List_HSTS.pdf
Attachment 3: BOM_List_HRTS.pdf
BOM_List_HRTS.pdf BOM_List_HRTS.pdf
  Draft   Fri Jun 7 07:32:00 2019 AlenaGeneral  
  147   Tue Jun 4 09:32:26 2019 Luis SanchezElectronicsCyMAC DOWNsCyMac's clock signal test

Luis, Rich, Rolf

With the help from Rolf a CyMac test was conducted. Apparently the clock signal on the blue CyMac box creates some sort of spikes or burts randomly. For this, we tried to measure or observed the clock signal and tried to catch this spikes effects.

1st. we verified if the signal still oscillating at the correct frequency, with oscilloscope we verified this and the signal appears behaving normally, 5vdc 65536Hz.

2nd. we tried a different approach by injecting a signal comming from a frequency generator into the adc or dac , this test has (for what we observed) same results as the setup with the Valon and buffer device (devices on blue box).

3rd. we observed during this test, that graphics appears to induce some burst or spikes, (not sure if this is real) these spikes were noticed after the display was awake from its screen saver process; after hitting the space bar, several spikes were noticed and then signal continues flat. Also, Rolf tried to open MatLab to see if this affect the process, in fact,  Matlab process also generates some spikes.

Attachment 1: CyMac_clock_test_060419.pdf
CyMac_clock_test_060419.pdf
  146   Tue Jun 4 08:22:17 2019 Luis SanchezElectronicsSatellite Box - OSEMLLO BOSEM TEST

Luis, Rich:

BOSEM's test was conducted; these units arrived from LLO (Stuart Aston) SN622 and SN224. A Set-up was prepared to conduct a test on clean environment. SN224 exhibited a nice signal performance, around 30µV/√Hz. This value is lower than values seen on our previous tests (60-70µV/√Hz). SN622 has a pronunced 1/f signal projection at lower frequencies and is more stable until it reach over 30Hz.

Attachment 1: LLO_OSEM_test.pdf
LLO_OSEM_test.pdf
  145   Wed May 29 10:39:50 2019 Luis SanchezElectronicsCLF RFPDS1800514 unit for modifications

Luis:

A CFL RFPD S1800514 was received from LLO for modifications and test. The unit tune-up follow E1900160 and T1200335.

The unit was modify by removing L5, L7, C35, L9, C38. Replacing U7 MAX4107 for LT6202. Placing jumpers at L5 and C38 and Replacing R22 20ohms for 49.9ohms and R26 2Kohms for 453ohms.

Simulation Results and calculations can be seen under S1800514.

Attachment 1: CFL_RFPD_Measured_Parameters_S1800514.pdf
CFL_RFPD_Measured_Parameters_S1800514.pdf CFL_RFPD_Measured_Parameters_S1800514.pdf
Attachment 2: Current_Noise_Calculated_Test_Results_from_S1800514.pdf
Current_Noise_Calculated_Test_Results_from_S1800514.pdf Current_Noise_Calculated_Test_Results_from_S1800514.pdf Current_Noise_Calculated_Test_Results_from_S1800514.pdf
  144   Tue May 28 08:02:01 2019 Luis SanchezElectronicsA-plusHSTS Altium wiring diagram

Luis:

I just finished an altium schematic from the HSTS filter, The dcc document is the number D1900169. The csv cable list file is also located in the same dcc number. Cables need a re-defined name or dcc part number, this will be generated later. A rack with possible U1 chassis are shown in the last page, overall I can said that this print will give us an idea on how to connect and how many chassis will be need for the final configuration. Also I need to verify if this arrange of chassis are correct.

Attachment 1: D1900169_Wiring_HSTS.pdf
D1900169_Wiring_HSTS.pdf D1900169_Wiring_HSTS.pdf D1900169_Wiring_HSTS.pdf
Attachment 2: D1900169_Wiring_HSTS_CableList.csv
Cable Number,Cable Info,Cable Length (Ft),From Description,From D-number,From Subsystem,From Rack,From U-height,From Connector Text,From Connector Location,From Connector Type, To Description,To D-number,To Subsystem,To Rack,To U-height,To Connector Text,To Connector Location,To Connector Type
VAC_TOP_3,Connector J4(C),4ft,6-BOSEM Array,D060218,SUS,SUS-R?,U?,TOP3,Front,D9F,Flange,D1000239,SUS,SUS-R?,U?,Q3,Rear,MicroD9
VAC_TOP_2,Connector J3(B),4ft,6-BOSEM Array,D060218,SUS,SUS-R?,U?,TOP2,Front,D9F,Flange,D1000239,SUS,SUS-R?,U?,Q2,Rear,MicroD9
VAC_TOP_1,Connector J2(A),4ft,6-BOSEM Array,D060218,SUS,SUS-R?,U?,TOP1,Front,D9F,Flange,D1000239,SUS,SUS-R?,U?,Q1,Rear,MicroD9
VAC_SIDE_6,Connector J3(B),4ft,6-BOSEM Array,D060218,SUS,SUS-R?,U?,SIDE,Front,D9F,Flange,D1000239,SUS,SUS-R?,U?,Q2,Rear,MicroD9
VAC_RIGHT_5,Connector J2(A),4ft,6-BOSEM Array,D060218,SUS,SUS-R?,U?,RIGTH,Front,D9F,Flange,D1000239,SUS,SUS-R?,U?,Q1,Rear,MicroD9
VAC_LEFT_4,Connector J5(D),4ft,6-BOSEM Array,D060218,SUS,SUS-R?,U?,LEFT,Front,D9F,Flange,D1000239,SUS,SUS-R?,U?,Q4,Rear,MicroD9
VAC_F2_8,Dxxxxxxx,4ft,Flange,D1000239,SUS,SUS-R?,U?,Flange,Front,D25M,Cable Bracket,D1001347,SUS,SUS-R?,U?,DeviceSide,Rear,D25
VAC_F1_7,Dxxxxxxx,4ft,Flange,D1000239,SUS,SUS-R?,U?,Flange,Front,D25M,Cable Bracket,D1001347,SUS,SUS-R?,U?,DeviceSide,Rear,D25
VAC_B2_10,Dxxxxxxx,15ft,Cable Bracket,D1001347,SUS,SUS-R?,U?,FlangeSide,Front,D25,D25 Vacuum Feedthrough,D?,SUS,SUS-R?,U?,DeviceSide,Rear,D25M
... 68 more lines ...
  143   Fri May 24 17:40:36 2019 Marie K. ProgressBS BRDsSecond BRD assembled

Today I built a second BRD mimicking the first one:

  • Bounce_brd2 = 2.269+2*1.025+screw = 4.936g
  • Roll_brd2 = 241+screw+spacer = 3.062g

I spent some time moving the masses along the blade axis to adjust the frequencies. As measured with the vibrometer, its characteristics are:

  • bounce mode: Q = 80 +/- 6  with f = 16.75 Hz (mean of 8 over 30 measurements).
  • roll mode: Q = 94 +/-7 with f = 24.25 Hz (mean of 8 over 30 measurements).

Then I installed the two BRDs on the suspension to get a rough estimation of the new BS Q. I had to double the drive (2.2V) to significantly excite the suspension. The data is noisy, the maximum of the frequency response is shifted compared to the excitation. The modes are hard to excite and damped very fast (figure 1 to 4) Ther might be an issue with the setup, I will have to chack it again without the BRDS.

I tried to scan the BS resonances with swept sin from the spectrum analyzer to finely measure the resonances of BS. However, the scans are not clean. The best scan I obtained is attached here (see figure 5). The resonances seem to be at 16.8 Hz and 24.4 Hz.

Attachment 1: BS_bounce_spectra.png
BS_bounce_spectra.png
Attachment 2: BS_bounce_ringdown.png
BS_bounce_ringdown.png
Attachment 3: BS_roll_spectra.png
BS_roll_spectra.png
Attachment 4: BS_roll_ringdown.png
BS_roll_ringdown.png
Attachment 5: best_TF_large_range.jpg
best_TF_large_range.jpg
  142   Thu May 23 10:07:04 2019 Luis SanchezElectronicsSatellite Box - OSEMBOSEM response as function of flag position

Luis:

I adapted a translation mechanism with a micrometer to measure BOSEM's flag position, the results can be seen on attached document.

Attachment 1: translation_mechanism_and_BOSEM_flag_response.pdf
translation_mechanism_and_BOSEM_flag_response.pdf
  141   Thu May 23 09:43:34 2019 Marie K.ProgressBS BRDsTuning of the BRD v4

Yesterday I spent some time adjusting the masses of the BRD to tune the resonances to the dummy BS resonances. Here is the summary of the attempts:

Masses/Freq Bounce - target = 16.70 Hz   Roll - target = 24.34 Hz  
Test 1 5.622g 14.9 Hz 2.620 g 27.9 Hz
Test 2

2.74+1.14+screw = 4.457 g

19 Hz 2.77+2*spacers+screw = 3.434 g 22.75 Hz
Test 3 2.74+1.14+screw+spacer = 4.497 g 19 Hz 2.77+spacer+screw = 3.389 g 22.5 Hz
Test 4 5.1 ; 2*1.02+2.49+screw = 5.167 g 16.25-16.5 Hz 2.54+screw = 3.16 g 23.75 Hz
Test 5 4.892; 2.27+1.02+0.98+screw = 4.896 g 17 Hz 3.083; 2.41+screw+spacer=3.063 g 24.5 Hz
Test 6 2.27+2*1.02+s=4.936 g 16.75 Hz  -  

Afterwards, I checked the values of the brd Q with the vibrometer:

  • Bounce mode: Q = 86 +/- 3 with f = 16.80 +/- 0.1 Hz (mean of 19 over 30 meausrements)
  • Roll mode: Q = 80 +/-2  with f = 24.6 +/ 0.15 Hz (mean of 16 over 30 measurements)

 

Then I installed the BRD on the dummy BS. The bounce mode Q is only reduced by a factor 2 (Q ~ 2600, see figure 1) but the roll mode is better damped (Q reduced to 600, see figure 2).

We need to remeasure the modes of the suspension alone and improve the tuning by further adjusting the masses.

Attachment 1: BS_bounce_damped.png
BS_bounce_damped.png
Attachment 2: BS_roll_damped.png
BS_roll_damped.png
  140   Tue May 21 16:37:06 2019 Marie K.ProgressBS BRDsFirst Q measurements

[Calum, Chub, Marie]

We installed a new support for the BOSEM (see first picture attached). The bosem translation stage seats on an aluminium plate, above a bridge in aluminium that is clamped to the BS cage.

Then we reconnected the satellite box in the setup, and we were able to restart the bosem excitation and the data acquisition. We removed the BRDs that had been previously installed on the pum. Then we measured the BS bounce and roll modes:

  • Excitation from the signal generator at 16.70 Hz. Resonance max is at 16.75 Hz. The measured Q is more than 4400 (see figure 2).
  • Excitation at 24.34 Hz. Resonance maximum is at 24.41 Hz. The measured Q is above 4300 (see figure 3).

We will scan around the resonances to check if the frequency is the sameas it was measured previously.

 

On the side, I measured the BRD that was assembled yesterday (brd_v4_n1) with the vibrometer:

  • The bounce frequency is 14.9 Hz, but the precision of the measurement is only 0.1 Hz (see for example figure 4). This is ~12% lower than expected. The mean value of 9 ringdown measurements indicate that the Q is about 85 +/- 3 (see an example figure 5).
  • The roll frequency is 27.9 Hz, with a precision of 0.1 Hz. This is ~ 14% higher than expected. The mean value of 5 ringdown measurements indicate that the Q is about 83 +/- 4. (see figure 6).

There is quite a large dispersion in the measurement of the Q, I have to improve (or pratice) the technique to excite only the main mode and avoid saturations (only 9 out of my 30 measurements are clean for bounce and 5 out of 30 for roll).

The Q's are lower than what had been measured with the version 3 of the blades, so we are probably going in the right direction (bounce mode Q ~ 135, roll mode Q~125 in T1700176-v7, section 6.2)

Attachment 1: Bosem_support.jpg
Bosem_support.jpg
Attachment 2: BS_bounce_ringdown.png
BS_bounce_ringdown.png
Attachment 3: BS_roll_ringdown.png
BS_roll_ringdown.png
Attachment 4: spectra_brdv4n1_bounce.png
spectra_brdv4n1_bounce.png
Attachment 5: ringdown_test6_Q90.png
ringdown_test6_Q90.png
Attachment 6: ringdown_roll_test4_Q80.png
ringdown_roll_test4_Q80.png
  139   Mon May 20 18:24:11 2019 Marie K. ProgressBS BRDsTest of the new springs

We received the 12 new springs made with a sandwich layer of Pyralux damping today (see dcc D1700188-v3). 

According to T1700176, the bounce and roll modes of the dummy BS in the modal lab are respectively 16.70 Hz and 24.34 Hz. Aiming for a Q~100 for the damped BS, the relative mass of the damper and the modes has to be > 2.0e-4. It means the masses have to be >= 5.556 g for the bounce mode and >= 2.580 g for the roll mode, taking into account the weight of the spring (0.168 g) and the screws (~0.49g each). I got: mb = 0.91 + 4.23 + screw = 5.622 g for bounce and mr = 2*1.06 + screw = 2.620 g for roll. 

Calum helped me taking the first measurement of the new BRD with the B&K software and the vibrometer. We found that the roll resonance is at 27.750 Hz (i.e. 14% too high), with a measured BRD Q of 59 (TBC).

Attachment 1: IMG_20190520_164859560.jpg
IMG_20190520_164859560.jpg
Attachment 2: IMG_20190520_184709517.jpg
IMG_20190520_184709517.jpg
  138   Mon May 20 14:05:56 2019 Marie K.ProgressBS BRDsElectronics moved to the Modal Lab

[Rich, Marie]

Today we moved the shelves containing the electronics for the BOSEM control and acquisition from Thomas Lab to the Modal Lab. The shelves are now near the dummy BS. We are planning to measure the bounce and roll modes of the BS suspension with and without BRDs. Please find attached some pictures for visual reference.

Luis fixed the satellite box for us, thanks! It is now in the lab, we still have to incorporate it back in the setup.

 

Attachment 1: IMG_20190520_101605561.jpg
IMG_20190520_101605561.jpg
Attachment 2: IMG_20190520_101623717.jpg
IMG_20190520_101623717.jpg
Attachment 3: IMG_20190520_101745757.jpg
IMG_20190520_101745757.jpg
Attachment 4: IMG_20190520_102700780.jpg
IMG_20190520_102700780.jpg
Attachment 5: IMG_20190520_141445408_HDR.jpg
IMG_20190520_141445408_HDR.jpg
  137   Thu May 9 13:43:47 2019 Craig WorleyHow ToModal TestingAdding Metadata title to B&K Graphs

To add the measurement name to the graph produced in the view tab, follow the following steps:

  1. Before taking the measurement, title the "Result Name" field the appropriate name for the experiment. Reference Picture 1 to see the location of this field. This will become the metadata that is used to identify series in the view tab.
  2. After the test has concluded, select into the Results mode (this button is seen in picture 1). In this tab, drag in the experiment results from the data tree into the results matrix. Once loaded, click the settings button (the wrench) that corresponds to the graph produced by the results matrix. Once this tab is opened up (it should look like the menu in picture 2), select the button that says "Click to add a title" located beside Legends. 
  3. A menu, shown in picture 3, will come up. On the top left window, select Results (Project->Device Under Test->Application->Test->Setup->Results). Now on the bottom left window select "Results Name". Using the arrows between the left and right side of the menu, add "Results Name" to the right side. This is the only thing needed in the menu. Close out of this window and proceed to the view tab.
  4. In the view tab, drag in the same data sets you entered into the results matrix in step 2 into the viewing area. This will display metadata titles like is seen in picture 4 at the top of the Figure. If you already entered the data into the viewing area before following the steps above, you need to re-enter the data into the viewing area. This just requires you to redrag the data sets into the graph. 
Attachment 1: Pic1.jpg
Pic1.jpg
Attachment 2: Pic2.jpg
Pic2.jpg
Attachment 3: Pic3.jpg
Pic3.jpg
Attachment 4: Pic4.jpg
Pic4.jpg
  135   Tue Apr 30 09:05:45 2019 Luis SanchezElectronicsA-plusWiring block diagram

Luis:

A block diagram for HSTS and HRTS were created, these might need a double check to verify if everything is well configured. Diagrams describe a connection between all components or chain of components for suspension HSTS or HRTS units. HSTS has a top, middle and bottom masses, while HRTS only has top and bottom masses. These diagrams will help when altium wiring diagram are created. Documents can be seen at E1900106 and E1900128.

Attachment 1: Visio-Suspension_Wiring_Block_Diagram_HSTS_E1900106v2.pdf
Visio-Suspension_Wiring_Block_Diagram_HSTS_E1900106v2.pdf Visio-Suspension_Wiring_Block_Diagram_HSTS_E1900106v2.pdf Visio-Suspension_Wiring_Block_Diagram_HSTS_E1900106v2.pdf Visio-Suspension_Wiring_Block_Diagram_HSTS_E1900106v2.pdf
Attachment 2: Suspension_Wiring_Block_Diagram_HRTS_E1900128-v1.pdf
Suspension_Wiring_Block_Diagram_HRTS_E1900128-v1.pdf Suspension_Wiring_Block_Diagram_HRTS_E1900128-v1.pdf Suspension_Wiring_Block_Diagram_HRTS_E1900128-v1.pdf Suspension_Wiring_Block_Diagram_HRTS_E1900128-v1.pdf
  134   Tue Apr 30 08:38:10 2019 Luis SanchezElectronicsCLF RFPDModify S1300534

Luis:

Unit S1300534 was modified, and we are using a toroid core inductor (T25-6, n=2.7) to generate a single resonant frequency of 6.25MHz. During modification process we noted that original photodiode had leads bended. Removed  photodiode is a PSL with number I>642G A 6106.  New photodiode was installed its number is C30642G 4778. Simulation and test results are very close, calculate current is 17.65x10-6A/Hz. Parameters of Tank Shunt Impedance are Rp=4.836Kohms, C=120.53pF(this include the variable cap and pcb traces) and L=5.3753x10-6H. MAX4107 was replaced by LT6202. During test we see Output Noise in dark of -131.8dBm/Hz (Simulation -131.95dBm/Hz), and Vrms of 57.47nVrms/√Hz (Simulation 56.47nVrms/Hz). See below calculations, test and simulation results. Documents can be also seen on S1300534.

 

Attachment 1: Current_Noise_Calculated_Test_Results_from_S1300534.pdf
Current_Noise_Calculated_Test_Results_from_S1300534.pdf Current_Noise_Calculated_Test_Results_from_S1300534.pdf Current_Noise_Calculated_Test_Results_from_S1300534.pdf
  133   Mon Apr 29 10:16:50 2019 Luis SanchezElectronicsPDHTrasnfer Function Board

Nick is troubleshooting Transfer Function Board from the PDH servo box, somehow the pcb board is getting hot and apparently pulling high current over some time. Nick isolated the input op amp THS4631 to check functionality, first ar dc then will do at ac. When testing at dc he found an oscillation, he pointed out the -12v power rail. He apply a DC voltage of 1v to the input and check the op amp, that in theory will give over -3.93v dc, Nick's reads -4.12v dc.

The op amp is getting hot, it might need a pcb thermal paddle to suppres the heat. Seems to me that the op-amp is having a hard time to drive the input capacitance. Nick will remove C20 from pcb and verify if this help with op amp heating and also with the oscillation seen at the negative voltage. More to come...

 

 

Attachment 1: Nick_troubleshooting_Transfer_Function_Board.pdf
Nick_troubleshooting_Transfer_Function_Board.pdf
Attachment 2: Re__Transfer_Function_PCB_Diagnosis.zip
  132   Wed Apr 24 14:31:57 2019 Luis SanchezElectronicsSatellite Box - OSEMLED current noise levels

Luis:

I am continue on testing the Satellite Box, after getting a dual low noise op amp ADA4898-2 and AD8599, I replace OP2177 and we observed a lower noise performance on LED current, the results can be seen below. Now the question is what would happen at the Photodiode site, for this I will take some new measurements after the IV preamplifier and compare if the reduction of LED Current noise has any effect on photodiode noise level.

In charts, device AD8599 in new configuration shows the best performance.

 

Attachment 1: AD8599_device_in_Channel_C_with_Sche.pdf
AD8599_device_in_Channel_C_with_Sche.pdf AD8599_device_in_Channel_C_with_Sche.pdf
  131   Wed Apr 17 12:52:33 2019 Luis SanchezElectronicsSatellite Box - OSEMCurrent Source test at lower frequency

Luis:

Results seen at previus eLog are not good, seem like SR785 used tool was not working correctly. Well a new set of plots were taking using a different SR785 tool, and the results can be seen below.  With this new set of plots we can now define if we really need filters located near voltage reference and TP106.   Also, I replot Tantalum 399-3766-2-ND dC/dT charcteristic. 

Test experiment were completed on 2 different channels, We compare results and both result channels appear to agreed.

Filter at the voltage reference appears to be working, limiting the noise coming out ADR421, but filter at current noise appears not to have any effect (TP106), do not know if this is limited by the instrument.

A simulation from  current source circuitry using low noise op amp and with some minor changes on resistor values delivers a lower noise performance. I will try this on the actual pcb board and see if any change really ocurrs.

Attachment 1: Satellite_Box_Channels_B_and_C_Test.pdf
Satellite_Box_Channels_B_and_C_Test.pdf Satellite_Box_Channels_B_and_C_Test.pdf Satellite_Box_Channels_B_and_C_Test.pdf Satellite_Box_Channels_B_and_C_Test.pdf Satellite_Box_Channels_B_and_C_Test.pdf Satellite_Box_Channels_B_and_C_Test.pdf Satellite_Box_Channels_B_and_C_Test.pdf Satellite_Box_Channels_B_and_C_Test.pdf
  130   Fri Apr 12 08:34:13 2019 Luis SanchezElectronicsSatellite Box - OSEMCurrent Source capacitors C113,C114,C115 and 10uF Tantalum test over temperature

Luis:

A test to verify filter bank capacitors (formed by C113(0.1uF), C114(OMIT), C115(100uF 20V)) on current source circuitry has any real effect on current apply to LED(OSEM). We took a differential voltage noise on R125(47ohms), and then convert this voltage to current. Results shows that at frequencies between 1-10Hz the filter plays a great roll, but not a frequencies >10Hz. For this test we conclude that we will keep this capacitors on the new design of  suspension satellite amplififer box fot A+.

Another observation is that a lower frequencies the current noise does not meet specifications from 1-5Hz, as seen on document "Sensors and Actuators for the Advanced Ligo Mirror Suspension" by L. Carbone et al. (Fig.8 page 9). We need to point out that this result is only for one unit been tested, not as the results from document (test more than 200 units) , and also we think that this reading is been limited by the SR785 used during test..

We also started to test a 10uF tantalum capacitor over temperature, tricky test since we want to keep a capacitor on a steady temperature while measuring the capacitance. For this initial test we managed to get some values from 30, 40, 50,60 Celcius. The inital capacitance value increase by a less than 1% over this range of temperature; seems that the reported documentation is correct or at least follow the same trend. Tantalum Capacitance increase when temperature increases.

Attachment 1: Satellite_Box_LED_Current_Supply_Test.pdf
Satellite_Box_LED_Current_Supply_Test.pdf Satellite_Box_LED_Current_Supply_Test.pdf Satellite_Box_LED_Current_Supply_Test.pdf
Attachment 2: Sensors_and_Actuators_for_the_Advanced_LIGO_Mirror_Suspensions_page_9.pdf
Sensors_and_Actuators_for_the_Advanced_LIGO_Mirror_Suspensions_page_9.pdf
  129   Tue Apr 9 09:50:58 2019 Luis SanchezGeneralHSTS wiringblock pieces for HSTS wiring and preliminary block diagram

Luis:

I started to create a block pieces that identify the units to be use during the HSTS wiring diagram, I started with, what I believe are the essential pieces, in these blocks I indicate what type of inputs and outputs for easy recognition when I create the actual library in Altium.  Also a preliminary block diagram was created for the HSTS chassis and wiring connection, after I verify this I will load the file on E1900106.

Attachment 1: Suspension_notes_blocks_diagram.pdf
Suspension_notes_blocks_diagram.pdf
Attachment 2: Visio-Suspension_Wiring_Block_Diagram_HSTS_E1900106.pdf
Visio-Suspension_Wiring_Block_Diagram_HSTS_E1900106.pdf Visio-Suspension_Wiring_Block_Diagram_HSTS_E1900106.pdf Visio-Suspension_Wiring_Block_Diagram_HSTS_E1900106.pdf
  128   Tue Apr 9 09:07:46 2019 Luis SanchezElectronicsPDHPDH system at MIT

Nick:

Finalize the assemble of PDH unit at MIT, unit appears to be functional, LCD shows intro banner when first start. Now I believe He will start testing PDH box with the help of Peter.

Attachment 1: Nick_PDH.pdf
Nick_PDH.pdf
Attachment 2: Nick_email_PDH.pdf
Nick_email_PDH.pdf
  127   Tue Mar 26 12:39:03 2019 Luis SanchezGeneralA-plusSuspension Satellite Amplifier panels

Luis:

Since we will have a review for the repackaging of UK Satellite Modules for A-plus, see E1900084. We created the front (D1900090) and rear (D1900091) panels for the Suspension Satellite Amplifier prototype, this was completed on the effort of having a visual idea on how the new chassis will look; it is important to have in mind that the configuration of the openings for the connectors (dSub connectors) and leds might change since we have not get any feedback from the Review Panel (the review will happen in the comming days or weeks). 

This panel will have a Green Quad Level LED Indicator with part number WP9345B / 4LGD for the visual indication of the 4 channels, if any channel fails the led will stay on, indicanting that a failure ocurred (high current seen at current source section). The vacuum chamber connection (Female DB25 Vacuum Tank), the Monitor to ADC (Male DB25 Analog Rack) and the Male D15 Coil Driver Interface (Analog Rack) will be set at front panel and the Local diagnostics Female DB37 connector will be set on the rear panel. More comments will be added later as we do some more work for a-plus.

 

Attachment 1: SS_Amplifier_Front_and_Real_panels.pdf
SS_Amplifier_Front_and_Real_panels.pdf
  126   Mon Mar 25 12:47:29 2019 Luis SanchezElectronicsPDHDiplexer module D1800249, Shield module D1700131

Nick and  Luis:

Several weeks ago Nick Pelepchan. EE student intern at MIT, shared with me some results from his diplexer module he built based on the DCC document D1800249. This diplexer is part of the PDH box. At the attachment section we can see the results he got during testing, appears that the unit is working as intented. At the moment He is working on the effort of assembling MIT PDH servo box.

Also, a few weeks back Nick was having some issues with Shield board D1700131.  The assembly house installed a polarized capacitor in wrong position, and the board was not getting the ±10volts, these power rails feed the digital potentiometer. After few trials He managed to identify the problem and fix the issue, now he continues working with the rest of the PDH servo box assembly.

 

Attachment 1: Nick_Pelepchan_EE_student_intern_at_MIT_diplexer.pdf
Nick_Pelepchan_EE_student_intern_at_MIT_diplexer.pdf Nick_Pelepchan_EE_student_intern_at_MIT_diplexer.pdf Nick_Pelepchan_EE_student_intern_at_MIT_diplexer.pdf Nick_Pelepchan_EE_student_intern_at_MIT_diplexer.pdf
Attachment 2: Nick_Pelepchan_EE_student_intern_at_MIT_-_shield.pdf
Nick_Pelepchan_EE_student_intern_at_MIT_-_shield.pdf
Attachment 3: Diplexer_emails.zip
Attachment 4: Shield_emails.zip
  125   Mon Mar 25 11:28:47 2019 Luis SanchezGeneralBRSProcurement orders from BRS project

Luis:

Created a table with all the material requested to create the BRS Heater Drivers, this can be seen at DCC C1900087.

 

Attachment 1: BRS_purchases_and_material.pdf
BRS_purchases_and_material.pdf
  124   Wed Mar 13 12:57:36 2019 Luis SanchezElectronicsBRSNoise Testing

Luis:

During BRS Heater driver noise test, an small oscillation was noticed, this oscillation make the signal bouncing around and erratic. After some electrical probing, we observed a sawtooth signal, this signal was found at the negative power rail. The signal was also seen at the current monitor capacitor "C17"(de-coupling capacitor for the -15v) and test point "T10"(which is the current monitor output). I installed a 10x10-6 F capacitor at c29 and C33, and also a 6.8x106F capacitor in parallel to C17. After these changes the signal improved and we continue to take the readings.

One observation is that the noise is dependent from Marchand amplifier board, and during testing we observed that these boards behave some how different in terms of noise performance.

Attachment 1: Sawtooth_signal_on_BRS_Heater_driver_D160454.pdf
Sawtooth_signal_on_BRS_Heater_driver_D160454.pdf Sawtooth_signal_on_BRS_Heater_driver_D160454.pdf
  123   Wed Mar 6 08:56:28 2019 Luis SanchezElectronicsGeneralBRS and CFL RFPD for LLO

Luis:

CFL RFPD: A modified unit was created for LLO site, its serial number is S1900053; for this unit we replace R1 resistor from 100ohm to 10kohm, this modification affect the transimpedance of the CFL RFPD, now we will have the ability of measure in the microAmp range. Another modification is that this unit adds capacitor C26 to circuitry, this capacitor was added to the circuit by a jumper wire located in L5 place; this capacitance is parallel to the photodiode capacitance, meaning that it will slightly increase the capacitance of the shunt circuit. This capacitor will help us to set the resonance frequency very close to 6.25MHz.

As the LHO unit we create an inductor, we use the same core from micrometals T25-6, for S1900053 the inductance is 5.881x10-6H at 6.25MHz. The Q=29.884, and the noise seen on spectrum analyzer was -129.0dBm/Hz, with these parameters we find a C=109.9pF, Rp=6.916Kohms. The shot noise equivalent seen at bench is 16.45x10-6Amp/Hz. It might be lower at the site on a better control environment.

BRS: Starting electrical test of BRS Heater Driver units, these test include the measurement of output voltages at different input levels and of course taking the transfer fucntion and noise spectrum from the chassis. For these tests, we use DB9 ADC terminal to measure the output voltage, while the DB9 DAC is use as source(pins 3-8 for voltage, pins 1-6 for ac signal). When measure the transfer function the relay jumper needs to be located on pin 2-3. For normal operation it has to be set on pins 1-2.

Attachment 1: CFL_RFPD_S1900053.jpg
CFL_RFPD_S1900053.jpg
Attachment 2: S1900053_charts.pdf
S1900053_charts.pdf
Attachment 3: Idc_Calculated_Test_Results_from_S1900053.pdf
Idc_Calculated_Test_Results_from_S1900053.pdf Idc_Calculated_Test_Results_from_S1900053.pdf Idc_Calculated_Test_Results_from_S1900053.pdf
Attachment 4: RFPD_MODIFICATION.jpg
RFPD_MODIFICATION.jpg
Attachment 5: assemble_units.PNG
assemble_units.PNG
Attachment 6: chassis_assembly.jpg
chassis_assembly.jpg
  122   Wed Feb 27 14:04:30 2019 Luis SanchezElectronicsCLF RFPDShipped CLF RFPD unit with number S1900052 to LHO

Luis:

Finally I got an inductor that give a resonant frequency of 6.25MHz, as describe before the core is a micro-metals T25-6 and the wire is a 34AWG copper magnetic. During testing we find out the the best device or the unit the has a better shot noise performance at 6.25MHz is the unit LT6202.  Using device LT6202 we saw a current noise of 20.77x10-6A/Hz, this value is for unit S1900052(which at parameters are C=108pF, Rp=6.298Kohms, Rs=8.75ohms,L=5.976x10-6H, Q=26.824, Fc=6.2534MHz). The measure noise of unit S1900052 is -128.8dBm/Hz, which gives a Vrms noise of 81.1867nVrms/sqrt(Hz). This unit will have a Tank Shunt Impedance with the next parameters L=5.976x10-6H, C=108pF, Rp=6.298Kohm, Gain=5. The simulation using these parameters with the addition of a differential capacitance of 3.2pF at the input of op amp gives a output voltage noise of 85.54x10-9V/sqrt(Hz), this value can also be read as -128.34dBm/Hz. For details on the calculations completed, please see document T1900075.

Unit S1900052 was shipped to LHO for testing on 02-26-19 and will arrive tomorrow 02-27-19. Now I need to continue on modifying the unit for LLO.

Attachment 1: charts.pdf
charts.pdf
Attachment 2: S1900052_calculations_&_simulation_results.pdf
S1900052_calculations_&_simulation_results.pdf S1900052_calculations_&_simulation_results.pdf
Attachment 3: CFL_RFPD_S1900052.jpg
CFL_RFPD_S1900052.jpg
  121   Tue Feb 26 11:04:27 2019 StephenProgressModal TestingOptical Cavity Eddy Current Damping

Looking at this work, the next step will be to re-generate plots with (1) improved resolution [Q should be resolved with (3dB width) > (10*resolution), if possible!] and (2) on longitudinal scales (either physical units or dB) to enable qualitative visual comparison. (also, I think the longitudinal movies you took didn't get posted?)

The following step would be to make a quantitative comparison of Q - see below table, intended as a mockup of the important information to summarize.

The key outcome of this effort is assessing whether the prototype damping scheme is working, and if so, what degrees of freedom should be incorporated into the final design of the damper. Hope this helps clarify why these requests are being made!

Comparison Table Mockup
Resonance Frequency (Hz) Prominent Direction (Long/Tran) Mode Shape Description Measurement technique Undamped 3 dB width (Hz) Undamped Q Best damped 3 dB width (Hz) Best damped Q

Best damping parameters

Resolution of FFT (Hz)
2.337 Transverse  Primary Pendulum Excitation, non-excitation .0163   .0241   1.5 mm spacing, 4 magnets, V .0125
2.438 or 2.425 Longitudinal Primary  Pendulum Excitation, non-excitation .0173   .0332   1.5 mm spacing, 4 magnets, V .0125
13.338 or w/e is next     Excitation, non-excitation (?)            
Quote:

Work in Progress

Contributors: Alexei, Stephen

Project: Optical Cavity Eddy Current Damping Mode Testing (BK Connect)

Summary: Non-excitation, transverse, and longitudinal excitation measurements by laser vibrometer (OMC and Ometron).  Recordings taken and fitted with fft spectrum to compare damping effects of different magnet configurations used for eddy current damping.

Description

The following measurements have been made:

  • No-excitation measurement in the transverse/longitudinal direction
  • Longitudinal excitation measurement undamped / 1.5mm damping / 3mm damping / 2 magnet centerline damping
  • Transverse excitation measurement undamped / 1.5mm damping / 3mm damping / 2 magnet centerline damping
  • Transverse/Longitudinal huddle tests

The pendulum mode of the optical cavity assembly was recorded when the assembly was undamped, damped by 4 magnets oriented at 45 degrees from the horizontal and with 2 magnets oriented along the centerline.  Distance of the magnets from the copper sheeting of the cavity was also tested (magnets kept at distance of 1.5mm and 3mm).

Note: OMC laser vibrometer has issue of signal LED having gone out. Prompted huddle tests.

Next Steps

  1. Alexei/Stephen - Finalize documentation from attached plots.  Calculate and compare Q's from fft spectra.

Photos/Figures

See attached

 

  120   Mon Feb 25 16:53:21 2019 AlexeiProgressModal TestingOptical Cavity Eddy Current Damping

Work in Progress

Contributors: Alexei, Stephen

Project: Optical Cavity Eddy Current Damping Mode Testing (BK Connect)

Summary: Non-excitation, transverse, and longitudinal excitation measurements by laser vibrometer (OMC and Ometron).  Recordings taken and fitted with fft spectrum to compare damping effects of different magnet configurations used for eddy current damping.

Description

The following measurements have been made:

  • No-excitation measurement in the transverse/longitudinal direction
  • Longitudinal excitation measurement undamped / 1.5mm damping / 3mm damping / 2 magnet centerline damping
  • Transverse excitation measurement undamped / 1.5mm damping / 3mm damping / 2 magnet centerline damping
  • Transverse/Longitudinal huddle tests

The pendulum mode of the optical cavity assembly was recorded when the assembly was undamped, damped by 4 magnets oriented at 45 degrees from the horizontal and with 2 magnets oriented along the centerline.  Distance of the magnets from the copper sheeting of the cavity was also tested (magnets kept at distance of 1.5mm and 3mm).

Note: OMC laser vibrometer has issue of signal LED having gone out. Prompted huddle tests.

Next Steps

  1. Alexei/Stephen - Finalize documentation from attached plots.  Calculate and compare Q's from fft spectra.

Photos/Figures

See attached

Attachment 1: Optical_Cavity.JPG
Optical_Cavity.JPG
Attachment 2: Optical_Cavity_mock_assembly.JPG
Optical_Cavity_mock_assembly.JPG
Attachment 3: Lab_setup_1.JPG
Lab_setup_1.JPG
Attachment 4: 4_magnet_damping_assembly.JPG
4_magnet_damping_assembly.JPG
Attachment 5: 4_magnet_side_view.JPG
4_magnet_side_view.JPG
Attachment 6: 2_magnet_damping_assembly.JPG
2_magnet_damping_assembly.JPG
Attachment 7: 2_magnet_side_view.JPG
2_magnet_side_view.JPG
Attachment 8: Transverse_undamped.MOV
Attachment 9: Transverse_damped.MOV
Attachment 10: Optical_Cavity_longitudinal_no_excitation_fft_spectrum_low_frequency_peaks.JPG
Optical_Cavity_longitudinal_no_excitation_fft_spectrum_low_frequency_peaks.JPG
Attachment 11: Optical_Cavity_longitudinal_no_excitation_fft_spectrum_higher_frequency_peak.JPG
Optical_Cavity_longitudinal_no_excitation_fft_spectrum_higher_frequency_peak.JPG
Attachment 12: Optical_Cavity_transverse_excitation_fft_spectrum.JPG
Optical_Cavity_transverse_excitation_fft_spectrum.JPG
Attachment 13: Optical_Cavity_longitudinal_excitation_fft_spectrum_peak_2.JPG
Optical_Cavity_longitudinal_excitation_fft_spectrum_peak_2.JPG
Attachment 14: Damped_2_magnets_centerline_comparison_longitudinal_excitation.JPG
Damped_2_magnets_centerline_comparison_longitudinal_excitation.JPG
Attachment 15: Damped_2_magnets_centerline_comparison_transverse_excitation.JPG
Damped_2_magnets_centerline_comparison_transverse_excitation.JPG
  119   Fri Feb 22 12:30:36 2019 Luis SanchezElectronicsBRSHeater Driver pcb boards

Luis:

The heater driver boards came back from stuffing, screaming circuit people could not stuffed the heat sink from power regulators, after some inspection was clear that a missing hole is the issue. I verify the gerber files I generated (gerber files appear OK) and asked Sunstone of what can be the issue. They email me back saying that Sunstone forgot to drill a hole on pcb boards; Dave from Sunstone said that after some investigation apparently they have some kind of problem with the CAM machine.  This drill holes are no important for the performance of the design, these are just holding holes for the heat sink, but since we wanted to know what was the problem I contacted them.

For the priority of the job I took the decision of stuff the boards, without inspecting the pcb boards, thinking that they will be manufacture as in the past. For now on I will verify that everything is ok before send out for stuffing o assembly.

Sunstone will send a new pcb boards for any future need heater driver.

 

Attachment 1: D1600454_75-S412097_W-10063091.jpg
D1600454_75-S412097_W-10063091.jpg
Attachment 2: sustone_email.pdf
sustone_email.pdf sustone_email.pdf sustone_email.pdf sustone_email.pdf
  118   Tue Feb 19 14:16:26 2019 Luis SanchezElectronicsCLF RFPDRFPD , BRS and Satellite Box

Todd, Luis on BRS:

With Todd's help we started to assemble some BRS chassis, and also continue on creating cable assemblis for these.

Luis on RFPD:

Rich and I created a table with some simulated results to compare with the bench test results. The noise seen on bench when testing THS4631 does not correlated to the noise seen on simulation. After some simulation analysis, we see a capacitance around of 8.9pF between the op amp inputs, I believe this capacitance is related to THS4631 data sheet differential input capacitance 3.9pF plus some more capacitance added from the copper traces. If we set the 8.9pF at the op amp inputs the onoise  seen is very close to the one observed at bench. I will continue testing with the new devices we ordered. LT6234, LT6202, OPA657 and AD829. For table seen below MAX4107 is the best device and is  delivering a better noise signal -127dBm/Hz (96nV/√Hz) at 6.235MHz. OPA657 is the next, but I have not test this to verified its behavior at 6.25MHz, need to determine capacitance and series resistance from the new devices. Waiting for these components that will arrive shortly.

Luis on Satellite Box:

Editing the old list of changes to be perform on the redesing of Satellite Box.

Attachment 1: BRS_assembly.pdf
BRS_assembly.pdf
Attachment 2: RFPD.pdf
RFPD.pdf
Attachment 3: D0901284_Suspension_UK_Satellite_Amplifier_Box_possible_changes_list.pdf
D0901284_Suspension_UK_Satellite_Amplifier_Box_possible_changes_list.pdf
  117   Thu Feb 14 15:48:26 2019 CraigProgressModal TestingVMD Review Questions

The following questions I came up with after listening in to the VMD design review on 2/13/2019

1) One method of tightening the VMD assembly is the allen key pattern in the top of the copper rod. If this process of tightening the assembly needs to be repeatable, how will the distortion of this pattern be taken into account? Just in the lab, the during re-assembly there has become noticeable distortion of this key pattern making it difficult to obtain the correct torque when tightening the assembly. 

2) I believe I heard there can be a +/- 1 Hz difference in resonance frequencies between Air/Vacuum, but how will epoxy react to this difference? Also, is there concerns of outgassing of the epoxy?

I hope to have those questions clarified by Stephen/Calum next week

  116   Tue Feb 12 13:55:32 2019 ElectronicsGeneralBRSHeater Driver Block Diagram

Luis:

A block diagram was created to help on the assembly of heater driver (BRS units), attached copy of document E1900039.

Attachment 1: Visio-Heater_Driver_E1900039.pdf
Visio-Heater_Driver_E1900039.pdf Visio-Heater_Driver_E1900039.pdf Visio-Heater_Driver_E1900039.pdf Visio-Heater_Driver_E1900039.pdf
  115   Tue Feb 12 13:40:33 2019 Luis SanchezElectronicsCLF RFPDSimulation and Bench Test Results

Luis:

Continue testing this RFPD using device MAX4107 and THS4631. The Simulation and Bench results from the MAX4107 are very close, at dark. Unfortunatly when replace the op-amp, using the THS4631 the noise at the output is very large and does not agree with the simulation results, we think that the device is some how was damage during installation. The readings are shown on attached document.

Results:

    nV/√Hz dBm/Hz   dBm/Hz
    Simulation Simulation   Agilent
MAX4107   90.4 -127.9   -125.8
THS4631   60.5 -131.35   -111
Attachment 1: RFPD_simulation_results.pdf
RFPD_simulation_results.pdf RFPD_simulation_results.pdf
  114   Wed Feb 6 13:53:30 2019 Luis SanchezElectronicsCLF RFPDTesting Results

Luis:

Continue testing CLF RFPD using MAX4107 and THS4631 devices as transimpedance amplifiers. An important observation is that the input capacitance from devices create different response over frequency. This was tested using the created toroid inductor with 6.05µH, CLF RFPD with MAX4107 has a resonance frequency of  6.24MHz, very close to the required 6.25MHz. Device THS4631 resonance frequency was noticed at 6.14MHz with 6.05µH(same coil than MAX4107).

Also a photocurrent test without preamplifier was perform using the MAX4107 device and noise was captured on a excel chart; after plotting the data two regions can be seen, appears that the Resistance (impedance of the system) behave more linearly between 0.06mA - 10mA, making noise increments more linear. At lower photocurrent between 0.01mA - 0.06mA the noise level is barely affected.

 

Attachment 1: Photocurrent_Test_Results.pdf
Photocurrent_Test_Results.pdf Photocurrent_Test_Results.pdf Photocurrent_Test_Results.pdf Photocurrent_Test_Results.pdf
ELOG V3.1.3-